# On the Convergence of Federated Learning Algorithms without Data Similarity

### v1

@@ I Introduction
Federated learning has gained significant popularity as a framework for training cutting-edge machine learning models using vast amounts of data collected from numerous resource-constrained devices, such as phones, tablets, and IoT devices. This approach allows these devices to individually train models using their private datasets without compromising sensitive information, [1, 2]. One common implementation of federated learning is the server-worker architecture, where a server aggregates information from local workers to update model parameters, which are then broadcast back to the workers. However, designing effective federated learning methods faces challenges such as dealing with high degrees of systems and statistical heterogeneity, as well as managing communication costs [1, 3, 4].
A popular approach in federated learning involves designing stochastic and distributed optimization algorithms that facilitate local updating. One such algorithm is FedAvg [1], also known as Local SGD [5, 6, 7, 8, 9, 10], which draws inspiration from stochastic gradient descent [11, 12, 13]. In this approach, each worker computes its local models based on a stochastic gradient using its private data and then communicates these models to the server responsible for updating the global models. Numerous other federated algorithms have emerged to amplify the training efficacy of FedAvg. For instance, FedProx [3, 14] and Proxskip [15] utilize proximal updates. Similarly, SCAFFOLD [16], FedSplit [17], and FedPD [18] harness variance reduction, operator splitting, and ADMM techniques respectively. FedPD was later refined into FedADMM [19] to expedite convergence.
The convergence behaviors of federated optimization algorithms in both homogeneous and heterogeneous data settings have been investigated in the literature. To model data heterogeneity, existing research often relies on data similarity assumptions, [3, 6, 14, 16, 18, 20]. One commonly used assumption measures the similarity between the local gradient at each worker and the global gradient. However, many existing works, e.g. [3, 16], require the step size to be tuned based on the data similarity to establish convergence. As a result, the step sizes tend to become extremely small, especially when the similarity between the local gradient at each worker and the global gradient is low. Moreover, these step sizes are generally impractical to compute since the data similarity is typically unknown in practice. In addition, many of the results only apply when the data similarity is high enough. Without these assumptions, the convergence of FedAvg for (strongly-) convex problems and FedADMM for non-convex problems are shown by [6] and [19], respectively. Their existing proof techniques and results cannot be applied to other federated learning algorithms, and they are limited to fixed step size strategies.
@@ I-A Contributions
The goal of this paper is to propose a unified framework for analyzing a broad family of federated learning algorithms for non-convex problems without data similarity assumptions. Our analysis is based on a general descent inequality that captures the convergence behaviors of several federated algorithms of interest. We derive novel sequence convergence theorems for three step size schedules commonly used in practice: fixed, diminishing, and step-decay step sizes. By applying these results, we establish convergence guarantees for popular federated algorithms. In particular, our convergence bound for FedAvg does not require restrictive assumptions unlike existing works in [21, 9, 22, 23], while our result for FedProx does not have the dependency of the step-size on the data similarity parameter in contrast to [3]. Finally, we demonstrate the effectiveness of these federated learning algorithms for deep neural network training over MNIST and FashionMNIST under different data similarity conditions.
@@ II Prior Works
II-A Data similarity assumptions.
@@
Two classical algorithms in federated learning include FedAvg [24] and FedProx [3, 14]. While FedAvg updates the global model by averaging local stochastic gradient descent updates, FedProx computes the model based on an average of local proximal updates. The convergence of both algorithms has been extensively analyzed under both homogeneous and heterogeneous data conditions. The data heterogeneity is often captured by different assumptions on data similarity which imply theoretical convergence performance of federated and decentralized algorithms, [7, 3, 25, 26, 27, 6, 5, 8, 10, 20]. For instance, FedAvg and FedProx are shown in [8] and [3], respectively, to converge slowly especially when the level of data similarity is low. While some limited works, e.g. [7, 6, 23, 19], have derived convergence results for federated learning algorithms without relying on data similarity assumptions, their results are confined to specific algorithms with fixed step sizes and cannot be extended to analyze other federated algorithms. Notably, our work makes a significant contribution by expanding the results in [7, 6], which only cover (strongly) convex problems, and by generalizing the results in [23], which requires restrictive assumptions on the Lipschitz continuity of the Hessian and on the bounded $4^{\text{th}}$-moment of the variance, i.e. $\mathbf{E}\|\nabla F_{i}(x;\xi)-\nabla f_{i}(x)\|^{4}\leq\sigma^{4}$ where $\nabla F_{i}(x;\xi)$ is the unbiased stochastic estimator of $\nabla f_{i}(x)$. In contrast, our results cover non-convex problems under standard assumptions, frequently encountered in the training of neural networks. In addition, we study various step size selection strategies, encompassing fixed, diminishing, and step-decay step sizes.
II-B Communication-efficient federated optimization.
@@
Communication bandwidth is a major performance bottleneck for federated algorithms [28]. This challenge becomes particularly pronounced in scenarios of high network latency, limited communication bandwidth, and when the communicated models are high dimensional. To alleviate the communication bottleneck, there are two common approaches. The first approach is to increase the number of local updates to reduce the number of communication rounds, often at the price of slow convergence speed, [1, 29]. The second approach is to reduce the number of communicated bits by applying compression, [28, 30, 31, 32, 33, 34, 35]. Compression can be sparsification (which keeps a few important vector elements) and/or quantization (which maps each vector element with infinite values into a smaller set of finite values). To further improve solution accuracy of algorithms using compression while saving communicated bits, error feedback mechanisms [36, 37, 38, 39, 40, 41, 42] and their variants, e.g. EF21 [43, 44, 45], have been extensively studied. The benefits of utilizing these approaches in federated learning have been explored by several works, e.g. [46, 47, 48]. Unlike these prior works, our framework can establish the convergence of error-feedback federated algorithms without data similarity assumptions. Our results also apply for stochastic, non-convex optimization unlike [46, 47], and do not assume bounded gradient-norm conditions unlike [48].
II-C step size schedules for stochastic optimization.
@@
Tuning step sizes is crucial to optimize the convergence performance of stochastic optimization algorithms. Using fixed step sizes for stochastic optimization algorithms guarantees the convergence towards the solution with the residual error, [49, 13]. To ensure the convergence of these algorithms towards the exact optimal solution, two common approaches are to use diminishing step sizes [50, 51, 52, 53, 54, 55], or step-decay step sizes [56, 57]. More recently, several works [58, 59, 60, 61] have proposed strategies for adjusting step sizes automatically to maximize the performance. However, theoretical convergence behaviors under different step size schedules are underexplored for federated learning. In this work, we unify the convergence of popular federated algorithms without data similarity assumptions when the step sizes are fixed, diminishing, and step-decay.
@@ IV Applications in Federated Learning
From Assumption 1 and by Cauchy-Schwartz’s inequality, the whole objective function $f(x)$ in (2) also has $L$-Lipschitz continuous gradient. Furthermore, the following inequalities are direct consequences from Assumption 1 [63, 64, 65]:
@@
Assumption 2 is standard to analyze the convergence of optimization methods using stochastic gradients, [66, 67, 28, 68]. In addition, since $\nabla f_{i}(x)=\mathbb{E}_{\xi_{i}}\nabla F_{i}(x;\xi_{i})$, this assumption implies that
@@
Assumption 2 is thus more general than the bounded second moment assumption, i.e. $\mathbb{E}_{\xi}\|\nabla F_{i}(x;\xi_{i})\|^{2}\leq\sigma^{2}$, which is used to derive the convergence of federated learning algorithms, see, e.g., [48, 5, 38, 69]. Also, this assumption is more relaxed than the bounded $4^{\text{th}}$-moment of the variance, i.e. $\mathbb{E}_{\xi}\|\nabla F_{i}(x;\xi_{i})-\nabla f_{i}(x)\|^{4}\leq\sigma^{4}$ studied in [23], because $\mathbf{E}\|\nabla F_{i}(x;\xi_{i})-\nabla f_{i}(x)\|^{2}\leq\sqrt{\mathbf{E}\|\nabla F_{i}(x;\xi_{i})-\nabla f_{i}(x)\|^{4}}$.
@@ IV-A Full-precision Federated Learning Algorithms
The full-precision federated learning algorithms are described formally in Algorithm 1.
@@ IV-A1 FedAvg
[…]
@@
By Proposition 1 and Theorem 1, FedAvg attains the $\mathcal{O}(1/K^{1/2})$ convergence for non-convex problems when $\gamma^{k}=\alpha^{k}/T$ and $\alpha^{k}=(\sqrt{6}L)^{-1}/\sqrt{K}$. Our result for FedAvg with fixed step sizes does not require data similarity assumptions, provides a faster rate than [23], and also does not require additional assumptions that restrict problem classes, e.g. the PL condition [22], the bounded gradient-norm condition [21], the condition that all eigenvalues of the mixing matrix must be less than $1$ [9], and the $3^{\text{rd}}$-order smooothness on a function (or the Lipschitz continuity on the Hessian )[23]. These comparisons between our convergence theorems and existing works were summarized in Table I. Furthermore, FedAvg converges at the $\mathcal{O}(1/K^{1-\nu})$ rate when $\alpha^{k}=(\sqrt{6}L)^{-1}/(k+1)^{\nu}$ for $\nu\in(1/2,1)$ from Theorem 2, and at the $\mathcal{O}(1/\log_{\alpha}(K))$ rate when $\alpha^{k}=(\sqrt{6}L)^{-1}/\alpha^{\lfloor k/T\rfloor}$ for $\alpha>1$ and $T=2K/\log_{\alpha}K$ from Theorem 3.
@@ IV-B Error-feedback Federated Learning Algorithms
These algorithms are formally summarized in Algorithm 2. Note that Algorithm 2 recovers FedPAQ [68] when we let $e_{i}^{k}=0$ for all $i,k$ and $\mathcal{T}_{\gamma F}(x)=x-\gamma\nabla F(x)$, and becomes Algorithm 1 when we set $e_{i}^{k}=0$ for all $i,k$ and $Q(x)=x$
@@
From Assumption 3, $\alpha$ implies the precision of a contractive compressor. For extreme cases, $Q(v)$ becomes close to $v$ as $\alpha$ is close to one. Contractive compressors cover the ternary quantizer [71] with $\alpha=1/d$, the scaled sign quantizer [38] with $\alpha=1/d$, and the Top-$K$ sparisifier [71, 44, 38] with $\alpha=K/d$.
@@ IV-B1 Error-feedback FedAvg
[…]
@@
From Proposition 3 and Theorem 1, 2, and 3, error-feedback FedAvg enjoys the $\mathcal{O}(1/K^{1/2})$, $\mathcal{O}(1/K^{1-\nu})$ and $\mathcal{O}(\log_{\alpha}K/\sqrt{K})$ convergence without data similarity assumptions, respectively, when $\alpha^{k}$ is fixed, diminishing and step-decay. Our $\mathcal{O}(1/K^{1-\nu})$ rate with diminishing step sizes is stronger than the $\mathcal{O}(1/\ln(K))$ rate by [48, Theorem 2]. In addition, in contrast to [48, Theorem 1], our result with fixed step sizes does not assume that the second moment is bounded, which is more restrictive than Assumption 2.
@@ V Numerical Experiments
We finally evaluated the performance of four different federated learning algorithms using three step size strategies to train deep neural network models over the MNIST [73] and FashionMNIST [74] datasets under various data similarity conditions. Both datasets contain $60000$ training images and $10000$ test images. Each $28\times 28$ grayscale image of the MNIST and FashionMNIST datasets is, respectively, one out of ten handwritten digits and one out of ten distinct fashion items. In particular, we implemented FedAvg, FedProx, error-feedback FedAvg, and error-feedback FedProx to solve the convolutional neural network (CNN) model using PyTorch [75]. This architecture contains a CNN with two 5x5 convolution layers, in which the first layer and second layer have 20 channels, and 50 channels, respectively, and each is followed by a 2x2 max pooling and ReLU activation function. Then, a fully connected layer with 500 units, a ReLU activation function, and a final softmax output layer formed our selected architecture. The total number of trainable parameters of this model is thus 431,080. All the numerical experiments were implemented in Python 3.8.6 and conducted on a computing server equipped with an NVIDIA Tesla T4 GPU with 16GB RAM. All source codes required for conducting and analyzing the experiments are made available online [^2] https://github.com/AliBeikmohammadi/FedAlgo_WO_DataSim/.
@@ V-B Hyper-parameters
For all algorithms, we set the number of communication rounds at $K=400$, chose the mini-batch size at $64$, and initialized the neural network weights using the default random initialization routines of the PyTorch framework. We chose the number of local updates at $T=30$ for FedAvg and error-feedback FedAvg, the learning rate of the inner solver for proximal updates at $0.1$ for FedProx and error-feedback FedProx, and $k$ to be $1\%$ of the trainable parameters (i.e. $k=4310$) for the top-$k$ sparsifier for error-feedback FedAvg and error-feedback FedProx. Furthermore, we employed three step size strategies: a) fixed step size with $c=2$, b) diminishing step size with $c=0.8$ and $\nu=0.51$, and c) step-decay step size with $\gamma_{0}=0.8$, $\alpha=2$, and $T=50$. For fair empirical comparisons, we ran the experiments using five distinct random seeds for network initialization. Figures 1 and 2 plot the average and standard deviation of training loss and test accuracy from running the algorithms with fixed step sizes over MNIST and FashionMNIST, respectively. We included additional experiments from running FedAvg, FedProx, error-feedback FedAvg, and error-feedback FedAvg with diminishing and step-decay step sizes over the MNIST dataset and FashionMNIST datasets. Particularly, we reported the result from running the algorithms with diminishing step sizes in Figures 3 and 4, and with the step-decay step sizes in Figures 5 and 6.
It should be noted that state-of-the-art methods might achieve higher test accuracy by utilizing more complex models and extensively tuning hyperparameters. However, the model introduced, along with the specified settings for algorithms and step sizes, adequately serves our purpose, which is to evaluate the optimization methods in the presence of various data similarities rather than aiming for the highest possible accuracy on these tasks.

### v2

@@ I Introduction
Federated learning has gained significant popularity as a framework for training cutting-edge machine learning models using vast amounts of data collected from numerous resource-constrained devices, such as phones, tablets, and IoT devices. This approach allows these devices to individually train models using their private datasets without compromising sensitive information [1]. One common implementation of federated learning is the server-worker architecture, where a server aggregates information from local workers to update model parameters, which are then broadcast back to the workers. However, designing effective federated learning methods faces challenges such as dealing with high degrees of systems and statistical heterogeneity, as well as managing communication costs [2].
A popular approach in federated learning involves designing stochastic and distributed optimization algorithms that facilitate local updating. One such algorithm is FedAvg [1], also known as Local SGD [3, 4], which draws inspiration from stochastic gradient descent [5]. In this approach, each worker computes its local models based on a stochastic gradient using its private data and then communicates these models to the server responsible for updating the global models. Numerous other federated algorithms have emerged to amplify the training efficacy of FedAvg. For instance, FedProx [2, 6] and Proxskip [7] utilize proximal updates. Similarly, SCAFFOLD [8], FedSplit [9], and FedPD [10] harness variance reduction, operator splitting, and ADMM techniques respectively. FedPD was later refined into FedADMM [11] to expedite convergence.
The convergence behaviors of federated optimization algorithms in both homogeneous and heterogeneous data settings have been investigated in the literature. To model data heterogeneity, existing research often relies on data similarity assumptions [6, 10]. One commonly used assumption measures the similarity between the local gradient at each worker and the global gradient. However, many existing works, e.g. [2, 8], require the step size to be tuned based on the data similarity to establish convergence. As a result, the step sizes tend to become extremely small, especially when the similarity between the local gradient at each worker and the global gradient is low. Moreover, these step sizes are generally impractical to compute since the data similarity is typically unknown in practice. In addition, many of the results only apply when the data similarity is high enough. Without these assumptions, the convergence of FedAvg for (strongly-) convex problems and FedADMM for non-convex problems are shown by [3] and [11], respectively. Their existing proof techniques and results cannot be applied to other federated learning algorithms, and they are limited to fixed step size strategies.
@@ I-A Contributions
The goal of this paper is to propose a unified framework for analyzing a broad family of federated learning algorithms for non-convex problems without data similarity assumptions, as shown in our analysis workflow in Figure 1. Our analysis is based on a general descent inequality that captures the convergence behaviors of several federated algorithms of interest. We derive novel sequence convergence theorems for three step size schedules commonly used in practice: fixed, diminishing, and step-decay step sizes. By applying these results, we establish convergence guarantees for popular federated algorithms. In particular, our convergence bound for FedAvg does not require restrictive assumptions unlike existing works in [12, 13, 14, 15], while our result for FedProx does not have the dependency of the step-size on the data similarity parameter in contrast to [2]. Finally, we demonstrate the effectiveness of these federated learning algorithms for deep neural network training over MNIST and FashionMNIST under different data similarity conditions.
Figure 1: Visual workflow of our analysis.
@@ II Prior Works
II-A Data Similarity Assumptions
@@
Two classical algorithms in federated learning include FedAvg [16] and FedProx [2, 6]. While FedAvg updates the global model by averaging local stochastic gradient descent updates, FedProx computes the model based on an average of local proximal updates. The convergence of both algorithms has been extensively analyzed under both homogeneous and heterogeneous data conditions. The data heterogeneity is often captured by different assumptions on data similarity which imply theoretical convergence performance of federated and decentralized algorithms [2, 17]. For instance, FedAvg and FedProx are shown in [17] and [2], respectively, to converge slowly especially when the level of data similarity is low. While some limited works, e.g. [4, 3, 15], have derived convergence results for federated learning algorithms without relying on data similarity assumptions, their results are confined to specific algorithms with fixed step sizes and cannot be extended to analyze other federated algorithms. Notably, our work makes a significant contribution by expanding the results in [4, 3], which only cover (strongly) convex problems, and by generalizing the results in [15], which requires restrictive assumptions on the Lipschitz continuity of the Hessian and on the bounded $4^{\text{th}}$-moment of the variance, i.e. $\mathbf{E}\|\nabla F_{i}(x;\xi)-\nabla f_{i}(x)\|^{4}\leq\sigma^{4}$ where $\nabla F_{i}(x;\xi)$ is the unbiased stochastic estimator of $\nabla f_{i}(x)$. In contrast, our results cover non-convex problems under standard assumptions, frequently encountered in the training of neural networks. In addition, we study various step size selection strategies, encompassing fixed, diminishing, and step-decay step sizes.
II-B Communication-efficient Federated Optimization
@@
Communication bandwidth is a major performance bottleneck for federated algorithms [18]. This challenge becomes particularly pronounced in scenarios of high network latency, limited communication bandwidth, and when the communicated models are high dimensional. To alleviate the communication bottleneck, there are two common approaches. The first approach is to increase the number of local updates to reduce the number of communication rounds, often at the price of slow convergence speed [1]. The second approach is to reduce the number of communicated bits by applying compression [19, 18]. Compression can be sparsification (which keeps a few important vector elements) and/or quantization (which maps each vector element with infinite values into a smaller set of finite values). To further improve solution accuracy of algorithms using compression while saving communicated bits, error feedback mechanisms [20] and their variants, e.g. EF21 [21], have been extensively studied. The benefits of utilizing these approaches in federated learning have been explored by several works, e.g. [22, 23]. Unlike these prior works, our framework can establish the convergence of error-feedback federated algorithms without data similarity assumptions. Our results also apply for stochastic, non-convex optimization unlike [22], and do not assume bounded gradient-norm conditions unlike [23].
II-C Step Size Schedules for Stochastic Optimization
@@
Tuning step sizes is crucial to optimize the convergence performance of stochastic optimization algorithms. Using fixed step sizes for stochastic optimization algorithms guarantees the convergence towards the solution with the residual error, [24, 5]. To ensure the convergence of these algorithms towards the exact optimal solution, two common approaches are to use diminishing step sizes [25, 26], or step-decay step sizes [27, 28]. More recently, several works [29, 30] have proposed strategies for adjusting step sizes automatically to maximize the performance. However, theoretical convergence behaviors under different step size schedules are underexplored for federated learning. In this work, we unify the convergence of popular federated algorithms without data similarity assumptions when the step sizes are fixed, diminishing, and step-decay.
@@ IV Applications in Federated Learning
From Assumption 1 and by Cauchy-Schwartz’s inequality, the whole objective function $f(x)$ in (2) also has $L$-Lipschitz continuous gradient. Furthermore, the following inequalities are direct consequences from Assumption 1 [32, 33]:
@@
Assumption 2 is standard to analyze the convergence of optimization methods using stochastic gradients [18, 34]. In addition, since $\nabla f_{i}(x)=\mathbb{E}_{\xi_{i}}\nabla F_{i}(x;\xi_{i})$, this assumption implies that
@@
Assumption 2 is thus more general than the bounded second moment assumption, i.e. $\mathbb{E}_{\xi}\|\nabla F_{i}(x;\xi_{i})\|^{2}\leq\sigma^{2}$, which is used to derive the convergence of federated learning algorithms, see, e.g., [23, 35]. Also, this assumption is more relaxed than the bounded $4^{\text{th}}$-moment of the variance, i.e. $\mathbb{E}_{\xi}\|\nabla F_{i}(x;\xi_{i})-\nabla f_{i}(x)\|^{4}\leq\sigma^{4}$ studied in [15], because $\mathbf{E}\|\nabla F_{i}(x;\xi_{i})-\nabla f_{i}(x)\|^{2}\leq\sqrt{\mathbf{E}\|\nabla F_{i}(x;\xi_{i})-\nabla f_{i}(x)\|^{4}}$.
@@ IV-A Full-precision Federated Learning Algorithms
The full-precision federated learning algorithms are described formally in Algorithm 1 and its visual workflow is shown in Figure 2.
Figure 2: Visual workflow of the full-precision federated learning algorithms.
@@ IV-A1 FedAvg
$\gamma_{k}$ $=\alpha^{k},\quad b_{1}=\sqrt{6}L^{2}T,\quad b_{2}=1/2,\quad\text{and}$
@@
By Proposition 1 and Theorem 1, FedAvg attains the $\mathcal{O}(1/K^{1/2})$ convergence for non-convex problems when $\gamma^{k}=\alpha^{k}/T$ and $\alpha^{k}=(\sqrt{6}L)^{-1}/\sqrt{K}$. Our result for FedAvg with fixed step sizes does not require data similarity assumptions, provides a faster rate than [15], and also does not require additional assumptions that restrict problem classes, e.g. the PL condition [14], the bounded gradient-norm condition [12], the condition that all eigenvalues of the mixing matrix must be less than $1$ [13], and the $3^{\text{rd}}$-order smoothness on a function (or the Lipschitz continuity on the Hessian )[15]. These comparisons between our convergence theorems and existing works were summarized in Table I. Furthermore, FedAvg converges at the $\mathcal{O}(1/K^{1-\nu})$ rate when $\alpha^{k}=(\sqrt{6}L)^{-1}/(k+1)^{\nu}$ for $\nu\in(1/2,1)$ from Theorem 2, and at the $\mathcal{O}(1/\log_{\alpha}(K))$ rate when $\alpha^{k}=(\sqrt{6}L)^{-1}/\alpha^{\lfloor k/T\rfloor}$ for $\alpha>1$ and $T=2K/\log_{\alpha}K$ from Theorem 3.
@@ IV-B Error-feedback Federated Learning Algorithms
These algorithms are formally summarized in Algorithm 2, and its visual workflow is shown in Figure 3. Note that Algorithm 2 recovers FedPAQ [37] when we let $e_{i}^{k}=0$ for all $i,k$ and $\mathcal{T}_{\gamma F}(x)=x-\gamma\nabla F(x)$, and becomes Algorithm 1 when we set $e_{i}^{k}=0$ for all $i,k$ and $Q(x)=x$
Figure 3: Visual workflow of the error-feedback federated learning algorithms.
@@
From Assumption 3, $\alpha$ implies the precision of a contractive compressor. For extreme cases, $Q(v)$ becomes close to $v$ as $\alpha$ is close to one. Contractive compressors cover the ternary quantizer [38] with $\alpha=1/d$, the scaled sign quantizer [39] with $\alpha=1/d$, and the Top-$K$ sparisifier [38, 21] with $\alpha=K/d$.
@@ IV-B1 Error-feedback FedAvg
$b_{2}$ $=\frac{1}{4},\quad\text{and}\quad b_{3}=2L\tilde{C}_{2}\Delta^{\inf}+\tilde{C}_{3}\sigma^{2}.$
@@
From Proposition 3 and Theorem 1, 2, and 3, error-feedback FedAvg enjoys the $\mathcal{O}(1/K^{1/2})$, $\mathcal{O}(1/K^{1-\nu})$ and $\mathcal{O}(\log_{\alpha}K/\sqrt{K})$ convergence without data similarity assumptions, respectively, when $\alpha^{k}$ is fixed, diminishing and step-decay. Our $\mathcal{O}(1/K^{1-\nu})$ rate with diminishing step sizes is stronger than the $\mathcal{O}(1/\ln(K))$ rate by [23, Theorem 2]. In addition, in contrast to [23, Theorem 1], our result with fixed step sizes does not assume that the second moment is bounded, which is more restrictive than Assumption 2.
@@ V Numerical Experiments
We finally evaluated the performance of four different federated learning algorithms using three step size strategies to train deep neural network models over two distinct datasets: MNIST [41] and FashionMNIST [42], under various data similarity conditions. Although it is feasible to explore additional datasets, we chose to focus on these two while considering different data distributions described in the following section to better highlight our key findings. Both datasets contain $60000$ training images and $10000$ test images. Each $28\times 28$ grayscale image of the MNIST and FashionMNIST datasets is, respectively, one out of ten handwritten digits and one out of ten distinct fashion items. In particular, we implemented FedAvg, FedProx, error-feedback FedAvg, and error-feedback FedProx to solve the convolutional neural network (CNN) model using PyTorch [43]. This architecture contains a CNN with two 5x5 convolution layers, in which the first layer and second layer have 20 channels, and 50 channels, respectively, and each is followed by a 2x2 max pooling and ReLU activation function. Then, a fully connected layer with 500 units, a ReLU activation function, and a final softmax output layer formed our selected architecture. The total number of trainable parameters of this model is thus 431,080. All the numerical experiments were implemented in Python 3.8.6 and conducted on a computing server equipped with an NVIDIA Tesla T4 GPU with 16GB RAM. All source codes required for conducting and analyzing the experiments are made available online [^2] https://github.com/AliBeikmohammadi/FedAlgo_WO_DataSim/.
@@ V-B Hyper-parameters
For all algorithms, we set the number of communication rounds at $K=400$, chose the mini-batch size at $64$, and initialized the neural network weights using the default random initialization routines of the PyTorch framework. We chose the number of local updates at $T=30$ for FedAvg and error-feedback FedAvg, the learning rate of the inner solver for proximal updates at $0.1$ for FedProx and error-feedback FedProx, and $k$ to be $1\%$ of the trainable parameters (i.e. $k=4310$) for the top-$k$ sparsifier for error-feedback FedAvg and error-feedback FedProx. Furthermore, we employed three step size strategies: a) fixed step size with $c=2$, b) diminishing step size with $c=0.8$ and $\nu=0.51$, and c) step-decay step size with $\gamma_{0}=0.8$, $\alpha=2$, and $T=50$. For fair empirical comparisons, we ran the experiments using five distinct random seeds for network initialization. Figures 4 and 5 plot the average and standard deviation of training loss and test accuracy from running the algorithms with fixed step sizes over MNIST and FashionMNIST, respectively. We included additional experiments from running FedAvg, FedProx, error-feedback FedAvg, and error-feedback FedAvg with diminishing and step-decay step sizes over the MNIST dataset and FashionMNIST datasets. Particularly, we reported the result from running the algorithms with diminishing step sizes in Figures 6 and 7, and with the step-decay step sizes in Figures 8 and 9.
It is worth mentioning that state-of-the-art methods might achieve higher test accuracy by employing more complex models and extensively tuning hyperparameters. However, the models we introduced, along with the specified settings for algorithms and step sizes, adequately serve our purpose: evaluating the optimization methods in the presence of various data similarities rather than achieving the highest possible accuracy on these tasks.